2) y + 3 y= 1 - 9x? diferansiyel denkleminin Y(0)=0, y(0) = 1 başlangia kosullarini sağlayan çözümünü bulunuz. (15 p.) ini

Find the solution of the differential equation according to the initial conditions of y (0) = 0, y '(0) = 1
Answer #1


Sol 1-9x² Given differential equation y + 3y! Y(0) = 0 , O) = 1 above equation can be rewritten as (22+ 3D) y auxilary equa 7 3;D 20 ,3 9x + 6 6x=1 =} (- 32° +32+- 4 = -x+x* y = Ye + Yp y = Gt Geox x+28 - 2 giveny (0) = 0 y 3 = OL 0 >> Ci+ C2 = 0 y